OpenStudy (anonymous):

One's intelligence quotient, or IQ, varies directly as a person's mental age and inversely as that person's chronological age. A person with the mental age of 25 and a chronological age of 20 has an IQ of 125. What is the chronological age of a person with a mental age of 40 and an IQ of 80?

OpenStudy (gw2011):

The chronological age is 50. Here is how I got the answer: Let MA=Mental Age Let CA=Chronological Age (MA/CA)(100)=IQ (40/CA)=80 (40)(100)=80(CA) (40)(100)/80=CA 50=CA